A car travels 130 miles in 2 hours. At the same speed, how far will the car travel in 3 hours?
Answer by TimothyLamb(4379)

---
s = d / t
s = 130/2
---
d = s * t
d = 130/2 * 3
---
answer:
d = 195 miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php